OpenStudy (anonymous):

solve using the multiplication principle -18=19y? would my answer be -18/19 or -18y/19

OpenStudy (anonymous):

It would be -18/19

OpenStudy (anonymous):

y=-18/19 because you multiply by 1/19 both sides

OpenStudy (anonymous):

\[-18=19y\] \[-(1/19)*18=(1/19)19y\] \[-18/19=y\]